OpenStudy (anonymous):

A line has a y-intercept of 7 and a slope of 8/7. Find the equation of the line. A. y = 7x + 8/7 B. y = 7x - 7/8 C. y = 8/7x + 7 D. y = 7/8x - 7

OpenStudy (***[isuru]***):

y = mx + b

OpenStudy (anonymous):

c

OpenStudy (***[isuru]***):

it's the slandered form of an straight line where m ( coefficient of "x" ) = slope b = y intercept

OpenStudy (***[isuru]***):

plug the values u got in to that m = 8/7 b = 7

OpenStudy (***[isuru]***):

u will get y = (8/7)x + 7... just like sienna said :)